OpenStudy (anonymous):

hey tropical33 can you see if i did this right

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

7x+5=6x+13 and my answer is 1

OpenStudy (anonymous):

sorry, but x=8

OpenStudy (anonymous):

2x/5=4

OpenStudy (anonymous):

i got 8

OpenStudy (anonymous):

x=10...

OpenStudy (anonymous):

i have 2 more can you help me please

OpenStudy (anonymous):

sure, just post them.

OpenStudy (anonymous):

Jack is three times as old as Lil. Their combined age is 52. How old is each person

OpenStudy (anonymous):

J=3L J+L=52 4L=52 L=13 and J=39

OpenStudy (anonymous):

the last problem...... the quotient of a number and 6